# A Comprehensive Overview of Large Language Models

### 27 Dec 2023 v7

@@ III-F Augmented LLMs
LLMs are capable of learning from the examples concatenated with the input, known as context augmentation, in-context learning (ICL), or few-shot prompting. They show excellent generalization to unseen tasks with few-shot prompting, enabling LLMs to answer queries beyond the capacity acquired during training [6, 51]. These emergent abilities allow for adapting the model without fine-tuning - a costly process. Aside from this, hallucination, producing inaccurate, unsafe or factually incorrect responses, is common for LLMs, which is avoided by augmenting contextual data. While the user can provide in-context samples in the query [50, 32], here we specifically refer to the methods that access external storage programmatically, calling them augmented LLMs.
@@
III-F1 Retrieval Augmented LLMs
@@
III-F2 Tool Augmented LLMs
@@
Figure 14: Illustration of parameter-efficient fine-tuning paradigms, where $x$ is input and $h$ is hidden state, figure courtesy [38].
III-G Efficient LLMs
III-G1 Parameter Efficient Fine-Tuning
@@
Fine-tuning LLMs with tens or hundreds of billions of parameters, such as GPT-3 (175B), BLOOM (176B), MT-NLG (540B), etc., is hardware-extensive and time-consuming. To avoid complete model fine-tuning, numerous parameter-efficient fine-tuning (PEFT) [40, 273, 41, 38, 39] are trying to reach full model fine-tuning performance at reduced costs. PEFT outperforms low-resource tasks, achieves comparable performance on medium-resource tasks, and underperforms on high-resource tasks compared to full fine-tuning [274]. An overview of different PEFT approaches is shown in Figure 14.
@@
Adapter Tuning: Adds a few trainable parameters within the transformer block. The adapter layer is a sequence of feature downscaling, non-linearity, and upscaling [102]. Variants of adapter tuning inject adapter layers sequentially [102] and parallel [275], whereas the mixture of adapter (AdaMix) [276] employs multiple adapter modules in a single layer. AdaMix routes input instances randomly to one of the multiple downscale and upscale modules. The mixture of adapters is averaged out for inference to avoid additional latency. Low-Rank Adaptation (LoRA) [233] learns low-rank decomposed matrices to freeze original weights. Learned weights are fused with the original weights for inference, avoiding latency.
@@ III-G2 Quantization
LLMs require extensive computing and memory for inference. Deploying the GPT-3 175B model needs at least 5x80GB A100 GPUs and 350GB of memory to store in FP16 format [281]. Such demanding requirements for deploying LLMs make it harder for smaller organizations to utilize them. Model compression is an effective solution but comes at the cost of degrading performance, especially at large scales greater than 6B. These models exhibit very large magnitude outliers that do not exist in smaller models [282], making it challenging and requiring specialized methods for quantizing LLMs [281, 283].
@@
Post-Training Quantization: Minimal or no training is required in this quantization type without compromising the model’s performance. LLM-8-bit [282] uses full-precision matrix multiplication for weights associated with outlier features and 8-bit for remaining. The lower precision multiplication outputs are converted to FP-16 and concatenated with others. The quantized models have homogenous word embeddings, degrading their performance. To fix this, token-level knowledge distillation is employed in [284] along with independent quantization scaling factors for each module due to varying weight distribution. Feature distributions are asymmetric and appear in different channels; outlier suppresion+ [285] shifts and scales per-channel activation distributions for effective quantization. SmoothQuant [281] quantizes activations and weights to INT8 by smoothing activations and migrating the quantization difficulty toward weights. It multiplies the inverse of the smoothing factor with weights, which introduces a few outliers in the weights but is easier to quantify than unsmoothed activations. OPTQ [283] uses the optimal brain compression (OBC) [286] algorithm to quantize the model layer-by-layer and update weights to compensate for quantization error. To improve speed and performance, OPTQ updates weights in arbitrary order, employs lazy updates, and uses better Cholesky kernels. Outlier-aware weight quantization (OWQ) [287] uses the OPTQ algorithm for quantization but assigns higher precision to vulnerable weights, causing outliers and lower precision for others.
@@
Quantization-Aware Training: To compensate for the performance degradation, the quantized model is fine-tuned in quantization-aware training (QAT) [288, 289, 290]. Alphatuning quantizes the model using binary coding quantization (BCQ) [291] and fine-tunes only quantization scaling factors. This approach improves performance over parameter-efficient fine-tuning of a pre-trained model and then quantizing and fine-tuning. Similarly, parameter-efficient and quantization-aware adaptation (PEQA) [292] reduces the precision of fully-connected layers and fine-tunes only quantization scaling parameters. LLM-QAT [290] generates training data from the pre-trained network and trains a quantized student model with knowledge distillation. QLoRA [289] fine-tunes 4-bit quantized pre-trained LLM with LoRA [233] using 4-bit normal float, which shows better performance over 4-bit integer and float.
III-G3 Pruning
@@
Unstructured Pruning: This kind of pruning removes less important weights without maintaining any structure. Existing LLM pruning methods take advantage of the unique characteristics of LLMs, uncommon for smaller models, where a small subset of hidden states are activated with large magnitude [282]. Pruning by weights and activations (Wanda) [293] prunes weights in every row based on importance, calculated by multiplying the weights with the norm of input. The pruned model does not require fine-tuning, saving large models’ computational costs. Outlier weighed layerwise sparsity (OWL) [296] extends Wanda with non-uniform layer pruning. It shows that the number of outliers varies for different layers; therefore, the model should have variable pruning ratios for better performance for every layer. Contrastive pruning (CAP) [297] iteratively prunes the model by training the sparse model using contrastive loss between pre-trained, fine-tuning, and snapshots of previous sparse models to learn task-specific and task-agnostic knowledge.
@@
Structured Pruning: Here, the parameters are removed in groups, rows, columns, or matrices, which speeds up the inference because of effective NVIDIA tensor core utilization [293]. LLM-Pruner [294] employs a 3-stage structured pruning strategy, identifying the groups of hidden states causing each other to activate during forward-pass, keeping important groups and removing less important ones, and fine-tuning the pruned model with LoRA. Sparsity-induced mask learning (SIMPLE) [298] prunes the network using learnable masks. Similarly, another method prunes LLMs by learning masks and removing unimportant rank-1 components of the factorized weight matrix [295].
@@
IV Findings & Insights
Training a billion-scale model is difficult as compared to a smaller model. LLMs are prone to various instabilities during training, such as hardware failure and instability. Other than this, LLMs exhibit different behaviors such as emergent abilities, improved zero-shot, few-shot, and reasoning abilities. Researchers report these essential details in their papers for results reproduction and field progress. We identify critical information in Table I and II such as architecture, training strategies, and pipelines that improve LLMs’ performance or other abilities acquired because of changes mentioned in section III.
V Model Configurations
@@
VI Datasets and Evaluation
Generating training and evaluation datasets is expensive because of the large-scale data demand of LLMs. Hence, datasets for training and benchmarking these models are topics of key importance. In Fig. 15, we show the distribution of the existing datasets for various NLP tasks. We restrict our distribution to only the most important tasks in the literature by including tasks with at least 20 datasets. LLMs can directly benefit from these datasets for training and evaluation. A summary of the training and evaluation datasets commonly used by LLMs is provided next.
VI-A Training Datasets
@@
VI-B Evaluation Datasets and Tasks
@@
Figure 15: A distribution of datasets proposed for different NLP tasks. We include only the tasks for which at least 20 datasets have already been proposed.
@@ VI-B2 Language Understanding
VI-B3 Story Cloze and Sentence Completion
@@
VI-B4 Physical Knowledge and World Understanding
@@
VI-B5 Contextual Language Understanding
@@ VI-B8 Mathematical Reasoning
VI-B9 Problem Solving and Logical Reasoning
@@
VI-B10 Cross-Lingual Understanding
@@
Task Dataset/Benchmark Model Model Size N-Shots Score
Multi-Task BIG-bench (B) Chinchilla 70B 5-shot 65.1
Gopher 280B 5-shot 53.97
PaLM 540B 5-shot 53.7
MMLU (B) GPT-4 - 5-shot 86.4
Gemini Ultra 5-shot 83.7
Flan-PaLM-2${}_{(f)}$ Large 5-shot 81.2
Language Understanding SuperGLUE (B) ERNIE 3.0 12B - 90.6
PaLM${}_{(f)}$ 540B - 90.4
T5 11B - 88.9
@@
Story Comprehension and Generation HellaSwag GPT-4 - 10-shot 95.3
Gemini Ultra 10-shot 87.8
PaLM-2 Large one shot 86.8
StoryCloze GPT3 175B few shot 87.7
PaLM-2 Large one shot 87.4
OPT 175B - 79.82
@@
Physical Knowledge and World Understanding PIQA PaLM-2 Large one shot 85.0
LLaMa 65B zero shot 82.8
MT-NLG 530B zero shot 81.99
TriviaQA PaLM-2 Large one shot 86.1
LLaMA-2 70B one shot 85.0
PaLM 540B one shot 81.4
Contextual Language Understanding LAMBADA PaLM 540B few shot 89.7
MT-NLG 530B few shot 87.15
PaLM-2 Large one shot 86.9
Commonsense Reasoning WinoGrande GPT-4 - 5-shot 87.5
PaLM-2 Large one shot 83.0
PaLM 540B zero shot 81.1
SIQA LLaMA 65B zero shot 52.3
Chinchilla 70B zero shot 51.3
Gopher 280B zero shot 50.6
Reading Comprehension BoolQ PaLM${}_{(f)}$ 540B - 92.2
T5 11B - 91.2
PaLM-2 Large one shot 90.9
Truthfulness Truthful-QA LLaMA 65B - 57
@@
Mathematical Reasoning MATH Gemini Ultra 4-shot 53.2
PaLM-2 Large 4-shot 34.3
LLaMa-2 65B 4-shot 13.5
GSM8K GPT-4 - 5-shot 92.0
PaLM-2 Large 8-shot 80.7
U-PaLM 540B - 58.5
Problem Solving and Logical Reasoning HumanEval Gemini${}_{(f)}$ Ultra zero shot 74.4
GPT-4 - zero shot 67.0
Code Llama 34B zero shot 48.8
VII Applications

### 20 Feb 2024 v8

@@ III-D Augmented LLMs
LLMs are capable of learning from the examples concatenated with the input, known as context augmentation, in-context learning (ICL), or few-shot prompting. They show excellent generalization to unseen tasks with few-shot prompting, enabling LLMs to answer queries beyond the capacity acquired during training [6, 55]. These emergent abilities allow for adapting the model without fine-tuning - a costly process. Aside from this, hallucination, producing inaccurate, unsafe, or factually incorrect responses, is common for LLMs, which is avoided by augmenting contextual data. While the user can provide in-context samples in the query [54, 32], here we specifically refer to the methods that access external storage programmatically, calling them augmented LLMs.
@@
III-D1 Retrieval Augmented LLMs
@@
III-D2 Tool Augmented LLMs
@@
III-E LLMs-Powered Agents
AI agents are autonomous entities, capable of planning, decision-making, and performing actions to achieve complex goals. In the older days, AI agents were rule-based, designed for narrow tasks, and had limited capabilities, such as Clippy [216] and Deep Blue [217]. In contrast to this, the LLMs capabilities to respond to dynamic scenarios have made it possible to incorporate them in diverse applications, including LLMs-powered agents [213, 205], where LLMs behave as the brain of agents. LLMs have been incorporated in web agents [155, 156], coding agents [218], tool agents [27, 212], embodied agents [26], and conversational agents [184], without requiring fine-tuning or minimal in some cases. Below we summarize the research in LLMs-based autonomous agents. For a more detailed discussion, please refer to [219, 220].
LLMs Steering Autonomous Agents: LLMs are the cognitive controllers of the autonomous agents. They generate plans, reason about tasks, incorporate memory to complete tasks and adapt the outline depending on the feedback from the environment. Depending on the acquired capabilities of LLMs, many methods fine-tune, propose a better prompting approach, or utilize different modules to enhance agents’ performance. Modules and strategies employed in autonomous agents are briefly discussed below.
Planning and Reasoning: Completing a complex task requires human-like logical thinking, planning necessary steps, and reasoning current and future directions. Prompting methods like chain-of-thoughts [98], tree-of-thoughts [100], and self-consistency [99] are central to agents, eliciting LLMs to reason its actions and choose among different paths for task completion. When LLMs are prompted with task description and sequence of actions, they can accurately generate plan actions without any fine-tuning [221]. Reasoning via planning (RAP) [222] incorporates a re-purposed LLM as a world model to reason about future outcomes and explore alternative paths for task completion. Retroformer [223] uses a retrospective LLM to improve main LLM planning and reasoning capabilities by providing helpful task cues.
Feedback: LLMs in open-loop systems generate plans and assume that the agent will complete it successfully. However, the actual scenario is different with failures and variable responses from the environment. To correctly complete tasks, many methods use LLMs in closed-loop where the action response is provided as feedback to the LLMs to re-assess and update the plan as required [224, 225, 226, 184]. Another direction of research exploits LLMs as reward functions to train reinforcement learning (RL) policies instead of humans [227].
Memory: LLMs can learn from the context provided in the prompt. In addition to internal memory, various systems employ external memory to save the response history. Reflexion [184] maintains an episodic memory to use previous responses as feedback to improve future decision-making. Retroformer [223] improves its responses by employing short-term and long-term memory, where short-term memory contains recent responses and long-term memory keeps summarized failed attempts to add in the prompt as reflection.
Multi-Agents Systems: LLMs can play user-defined roles and behave like a specific domain expert. In multi-agent systems, each LLM is assigned a unique role, simulating human behavior and collaborating with other agents to complete a complex task [218, 228].
LLMs in Physical Environment: LLMs are good at instruction-following, however, utilizing them for physically grounded tasks requires adaptation, as they lack real-world knowledge. This could lead to generating illogical responses for a particular physical situation [229, 26]. SayCan [229] make LLMs aware of the available low-level task operations. LLM (Say) builds a high-level plan to complete the task and a learned affordance function (Can) explores the possibility of executing the plan in the real world. SayCan uses RL to train the language-conditioned affordance function. PaLM-E enables the LLM to solve grounded tasks by training multi-modal LLM feeding inputs directly from the sensors.
Manipulation: In the area of manipulation [225, 230], LLMs enhance a robot’s dexterity and adaptability, excelling in tasks like object recognition, grasping, and collaboration. They analyze visual and spatial information to determine the most effective approach to interact with objects.
Navigation: LLMs enhance a robot’s ability to navigate complex environments with precision and adaptability [231, 232, 233, 234]. They generate feasible paths and trajectories for robots, accounting for intricate environmental details [235]. This ability is valuable in scenarios requiring precise and dynamically adaptable navigation in environments like warehouses, transport, healthcare facilities, and residences.
@@ III-E LLMs-Powered Agents
Figure 14: Illustration of parameter-efficient fine-tuning paradigms, where $x$ is input and $h$ is hidden state, figure courtesy [38]. Parallel adapter and LoRA fall in the adapter tuning category
III-F Efficient LLMs
Deploying LLMs in production is expensive. Reducing their running costs while preserving performance is an appealing area of research. This section summarizes the approaches suggested to enhance LLMs efficiency.
III-F1 Parameter Efficient Fine-Tuning
@@
Fine-tuning LLMs with tens or hundreds of billions of parameters, such as GPT-3 (175B), BLOOM (176B), MT-NLG (540B), etc., is computationally-intensive and time-consuming. To avoid complete model fine-tuning, numerous parameter-efficient fine-tuning (PEFT) techniques [40, 236, 41, 38, 39] try to achieve acceptable model fine-tuning performance at reduced costs. As compared to full fine-tuning [237], PEFT performs better in low-resource setups, achieves comparable performance on medium-resource scenarios, and performs worse than full fine-tuning under high-resource availability. An overview of different PEFT approaches is shown in Figure 14.
@@
Adapter Tuning: Adds a few trainable parameters within the transformer block. The adapter layer is a sequence of feature downscaling, non-linearity, and upscaling [101]. Variants of adapter tuning inject adapter layers sequentially [101] and in parallel [38], whereas the mixture of adapter (AdaMix) [238] employs multiple adapter modules in a single layer. AdaMix routes input instances randomly to one of the multiple downscale and upscale modules. The mixture of adapters is averaged out for inference to avoid additional latency. Low-Rank Adaptation (LoRA) [239] learns low-rank decomposed matrices to freeze original weights. Learned weights are fused with the original weights for inference, avoiding latency.
@@ III-F2 Quantization
LLMs require extensive computing and memory for inference. Deploying 175B parameter GPT-3 model needs at least 5x80GB A100 GPUs and 350GB of memory to store in FP16 format [44]. Such demanding requirements for deploying LLMs make it harder for smaller organizations to utilize them. Model compression is an effective solution but comes at the cost of degrading performance, especially at large scales greater than 6B. These models exhibit very large magnitude outliers that do not exist in smaller models [244], making it challenging and requiring specialized methods for quantizing LLMs [44, 245].
@@
Post-Training Quantization: Minimal or no training is required in this type of quantization, without significantly compromising the model performance. LLM-8-bit [244] uses full-precision matrix multiplication for weights associated with outlier features and 8-bit for remaining. The lower precision multiplication outputs are converted to FP-16 and concatenated with others. The quantized models have homogenous word embeddings, which may degrade their performance. To fix this, token-level knowledge distillation is employed in [45] along with independent quantization scaling factors for each module due to varying weight distribution. Feature distributions are asymmetric and appear in different channels; outlier suppression [246] shifts and scales per-channel activation distributions for effective quantization. SmoothQuant [44] quantizes activations and weights to INT8 format by smoothing activations and migrating the quantization difficulty toward weights. It multiplies the inverse of the smoothing factor with weights, which introduces a few outliers in the weights but is easier to quantify than unsmoothed activations. OPTQ [245] uses the optimal brain compression (OBC) [247] algorithm to quantize the model layer-by-layer and update weights to compensate for quantization error. To improve speed and performance, OPTQ updates weights in arbitrary order, employs lazy updates, and uses better Cholesky kernels. Outlier-aware weight quantization (OWQ) [248] uses the OPTQ algorithm for quantization but assigns higher precision to vulnerable weights, causing outliers and lower precision for others.
@@
Quantization-Aware Training: To compensate for performance degradation, a quantized model is fine-tuned in quantization-aware training (QAT) [249, 250, 251]. Alphatuning quantizes the model using binary coding quantization (BCQ) [252] and fine-tunes only quantization scaling factors. This approach improves performance over parameter-efficient fine-tuning of a pre-trained model and then quantizing and fine-tuning. Similarly, parameter-efficient and quantization-aware adaptation (PEQA) [253] reduces the precision of fully-connected layers and fine-tunes only quantization scaling parameters. LLM-QAT [251] generates training data from the pre-trained network and trains a quantized student model with knowledge distillation. QLoRA [250] fine-tunes 4-bit quantized pre-trained LLM with LoRA [239] using 4-bit normal float, which shows better performance over 4-bit integer and float.
III-F3 Pruning
@@
Unstructured Pruning: This kind of pruning removes less important weights without maintaining any structure. Existing LLM pruning methods take advantage of the unique characteristics of LLMs, uncommon for smaller models, where a small subset of hidden states are activated with large magnitude [244]. Pruning by weights and activations (Wanda) [254] prunes weights in every row based on importance, calculated by multiplying the weights with the norm of input. The pruned model does not require fine-tuning, thereby saving computational costs. Outlier weighed layerwise sparsity (OWL) [256] extends Wanda with non-uniform layer pruning. It shows that the number of outliers varies for different layers; therefore, the model should have variable pruning ratios for better performance for every layer. Contrastive pruning (CAP) [43] iteratively prunes the model by training the sparse model using contrastive loss between pre-trained, fine-tuning, and snapshots of previous sparse models to learn task-specific and task-agnostic knowledge.
@@
Structured Pruning: Here, the parameters are removed in groups, rows, columns, or matrices, which speeds up the inference because of effective hardware tensor core utilization [254]. LLM-Pruner [42] employs a 3-stage structured pruning strategy, identifying the groups of hidden states causing each other to activate during forward-pass, keeping important groups and removing less important ones, and fine-tuning the pruned model with LoRA. Sparsity-induced mask learning (SIMPLE) [257] prunes the network using learnable masks. Similarly, another method prunes LLMs by learning masks and removing unimportant rank-1 components of the factorized weight matrix [255].
III-G Multimodal LLMs
III-H Summary and Discussion
III-H1 Architecture
Layer Normalization: is found to have a significant effect on the performance and training stability of LLMs. Pre-norm, that is normalizing inputs rather than outputs, is more common among LLMs stabilizing the training [6, 125, 103]. BLOOM [13] and AlexaTM [120] utilize an additional layer normalization before embedding layer to stabilize the training of large-scale models, while the model’s zero-shot generalization ability can be negatively impacted [13]. However, another study [33] finds that pre-norm degrades fine-tuned model performance as compared to post-norm, and there are no stability benefits of pre-norm beyond the 100B scale. Therefore, GLM-130B [33] used deep-norm which is a variant of post-norm for better downstream task performance after fine-tuning.
Positional Encoding: effect performance and training stability of LLMs like other building blocks of a model. BLOOM [13] finds ALiBi outperforming learned and rotary positional encodings. Contrary to this, GLM-130B [33] identifies rotary positional encoding better than ALiBi. So, there is no conclusion in literature about the positional encodings yet.
Parallel Attention: where attention and feed-forward layers are parallel to each other rather than sequential in transformer block has shown to reduce training time by 15%. There is no evidence of performance drop due to this change in literature and used by the models PaLM [15], GPT-NeoX [113], and CodeGen [129].
Mixture of Experts: allows easily scaling model to trillion of parameters [128, 116]. Only a few experts are activated during the computation making them compute-efficient. The performance of MoE models is better than the dense models for the same amount of data and requires less computation during fine-tuning to achieve performance similar to the dense models as discussed in [116]. MoE architectures are less prone to catastrophic forgetting, therefore are more suited for continual learning [128]. Extracting smaller sub-models for downstream tasks is possible without losing any performance, making MoE architecture hardware-friendly [128].
Sparse vs Dense Activated: GPT-3 [6] uses sparse transformers [67] whereas GLaM [116] and PanGu-$\sum$ [128] use MoE [117] architecture to lower computational costs and increase the model size and capacity. According to the literature, sparse modules do not degrade the model’s performance [67]. However, more experiments are required to verify this statement.
III-H2 Training Strategies
Training models at a huge scale require tricks to reduce training costs, avoid loss divergence and achieve better performance. We summarize and discuss some of these key tricks used in different LLMs.
Mixed Precision: is a famous method for LLMs to reduce memory usage and improve training efficiency. In mixed precision, forward and backward passes are performed in FP16 format whereas optimizer states and master weights are kept in FP32 format [115]. A drawback associated with this format change is training instability due to a smaller value range resulting in loss spikes [33]. An alternative to FP16 is BF16 which has a comparatively larger range and performs some precision-sensitive operations like gradient accumulation and softmax in FP32 [13]. BF16 has better performance and training stability but uses more memory and is supported on specific hardware, for example, A100 GPUs. Therefore, its adoption in LLMs is limited.
Training Instability: is a common issue in LLMs where loss divergence or spiking is observed multiple times during training. This happens in the presence of gradient clipping [15]. To mitigate this problem, many approaches suggest restarting training from an earlier checkpoint [15, 33, 116], skipping 200-500 earlier data batches at the point of divergence in [15] and re-shuffling batches in [116]. The embedding layer gradient shrink proves to further stabilize the training as its gradient norm is significantly larger than the other layers [33]. Another suggestion to improve training stability for larger models is not to use biases in dense and norm layers as in [15].
Weight Initialization: plays a significant role in model convergence and training stability. GPT-NeoX [113] initializes feed-forward layers before residuals with $\frac{2}{L\sqrt{d}}$ as in [142] and other layers with small initialization scheme [287]. This avoids activations growing exponentially with the increasing depth. MT-NLG [112] found higher variance for weight initialization leads to unstable training, hence validating small initialization scheme [287]. Various models perform random weight initialization which can cause bad initialization, Galactica [137] suggests a longer warmup to negate the effect.
Learning Rate: is important for stable training. It is suggested to use a lower value [13, 15, 122] with warmup and decay (cosine or linear). Usually, the learning rate is within the range $1e^{-4}$ to $8e^{-4}$. Moreover, MT-NLG (530B) [112] and GPT-NeoX (20B) [113] suggest interpolating learning rates based on the model size using the GPT-3 [6] models ranging between 13B and 175B. This avoids tuning the learning rate hyperparameter.
Training Parallelism: 3D parallelism, a combination of data, pipeline and tensor parallelism, is the most utilized training parallelism approach in LLMs [33, 15, 14, 13, 112, 110, 107]. In addition to the 3D parallelism, BLOOM [13] uses zero optimizer [37] to shard optimizer states. PanGu-$\alpha$ [103] and PanGu-$\Sigma$ [128] go beyond the 3D parallelism and apply 5D parallelism which additionally contains optimizer parallelism and rematerialization.
Mode Switching: adds task-related tokens at the beginning of the text during training. These tokens refer to the natural language understanding and natural language generation tasks which are shown to improve the downstream task performance in [123, 122, 120]. During fine-tuning and inference, tokens are appended based on the downstream tasks.
Controllable Text Generation: Generating credible and controlled text from a pre-trained model is challenging. GPT-3 [6] and other LLMs use in-context learning to control generated text. While in-context learning helps in controlling the generated text, ERNIE 3.0 Titan [35] suggests using adversarial loss to rank its generated text for credibility and soft prompts such as genre, topic, keywords, sentiment, and length for better control on generated text.
III-H3 Supervised Models vs Generalized Models
III-H4 Zero-Shot vs Few-Shot
III-H5 Encoder vs Decoder vs Encoder-Decoder
@@
IV Model Configurations
@@
V Datasets and Evaluation
Generating training and evaluation datasets is expensive because of the large-scale data demand of LLMs. Hence, datasets for training and benchmarking these models are topics of key importance. A summary of datasets commonly used by LLMs is provided next.
V-A Training Datasets
@@
V-B Evaluation Datasets and Tasks
@@ V-B2 Language Understanding
V-B3 Story Cloze and Sentence Completion
@@
V-B4 Physical Knowledge and World Understanding
@@
V-B5 Contextual Language Understanding
@@ V-B8 Mathematical Reasoning
V-B9 Problem Solving and Logical Reasoning
@@
V-B10 Cross-Lingual Understanding
@@
Task Dataset/Benchmark Top-1 Top-2 Top-3
Model (Size) Score (N-shots) Model (Size) Score (N-shots) Model (Size) Score (N-shots)
Multi-Task BIG-bench (B) Chinchilla (70B) 65.1 (5-shot) Gopher (280B) 53.97 (5-shot) PaLM (540B) 53.7 (5-shot)
MMLU (B) GPT-4 (-) 86.4 (5-shot) Gemini (Ultra) 83.7 (5-shot) Flan-PaLM-2${}_{(f)}$ (Large) 81.2 (5-shot)
Language Understanding SuperGLUE (B) ERNIE 3.0 (12B) 90.6 (-) PaLM${}_{(f)}$ (540B) 90.4 (-) T5 (11B) 88.9 (-)
@@
Story Comprehension and Generation HellaSwag GPT-4 (-) 95.3 (10-shot) Gemini (Ultra) 87.8 (10-shot) PaLM-2 (Large) 86.8 (one shot)
StoryCloze GPT3 (175B) 87.7 (few shot) PaLM-2 (Large) 87.4 (one shot) OPT (175B) 79.82 (-)
@@
Physical Knowledge and World Understanding PIQA PaLM-2 (Large) 85.0 (one shot) LLaMa (65B) 82.8 (zero shot) MT-NLG (530B) 81.99 (zero shot)
TriviaQA PaLM-2 (Large) 86.1 (one shot) LLaMA-2 (70B) 85.0 (one shot) PaLM (540B) 81.4 (one shot)
Contextual Language
Understanding
LAMBADA PaLM (540B) 89.7 (few shot) MT-NLG (530B) 87.15 (few shot) PaLM-2 (Large) 86.9 (one shot)
Commonsense Reasoning WinoGrande GPT-4 (-) 87.5 (5-shot) PaLM-2 (Large) 83.0 (one shot) PaLM (540B) 81.1 (zero shot)
SIQA LLaMA (65B) 52.3 (zero shot) Chinchilla (70B) 51.3 (zero shot) Gopher (280B) 50.6 (zero shot)
Reading Comprehension BoolQ PaLM${}_{(f)}$ (540B) 92.2 (-) T5 (11B) 91.2 (-) PaLM-2 (Large) 90.9 (one shot)
Truthfulness Truthful-QA LLaMA (65B) 57 (-)
@@
Mathematical Reasoning MATH Gemini (Ultra) 53.2 (4-shot) PaLM-2 (Large) 34.3 (4-shot) LLaMa-2 (65B) 13.5 (4-shot)
GSM8K GPT-4 (-) 92.0 (5-shot) PaLM-2 (Large) 80.7 (8-shot) U-PaLM (540B) 58.5 (-)
Problem Solving and
Logical Reasoning
HumanEval Gemini${}_{(f)}$ (Ultra) 74.4 (zero shot) GPT-4 (-) 67.0 (zero shot) Code Llama (34B) 48.8 (zero shot)
VI Applications